OpenStudy (anonymous):

dale tries to eat at least 5 servings of fruit each day. for the first six days she had 5,6,3,5,4 how many servings of fruit should dale eat on sat to average at least 5 servings per day for the week

jimthompson5910 (jim_thompson5910):

it says "for the first six days" but it lists only 5 numbers (5,6,3,5,4) is this a typo?

OpenStudy (anonymous):

yes it should have another 4 listed after the 4

jimthompson5910 (jim_thompson5910):

so the full list is 5,6,3,5,4,4 ??

OpenStudy (anonymous):

yes

jimthompson5910 (jim_thompson5910):

let x = amount eaten on saturday if you average the 7 days, youll get this A = (5+6+3+5+4+4+x)/7 where A = average amount of servings

jimthompson5910 (jim_thompson5910):

you want this average to be 5 or more, so plug in A = 5 to get 5 = (5+6+3+5+4+4+x)/7 now solve for x

OpenStudy (anonymous):

4

jimthompson5910 (jim_thompson5910):

it's too low

OpenStudy (anonymous):

im getting 27 +x/7 though

jimthompson5910 (jim_thompson5910):

ok so 5 = (5+6+3+5+4+4+x)/7 5 = (27+x)/7

jimthompson5910 (jim_thompson5910):

now multiply both sides by 7 7*5 = 27+x 27+x = 35 x = 35 - 27 x = 8

OpenStudy (anonymous):

ohhhhh ok 8!

jimthompson5910 (jim_thompson5910):

somehow you cut it in half i think

jimthompson5910 (jim_thompson5910):

but yes, x = 8

OpenStudy (anonymous):

so she would have to eat more than 8 to get the average

jimthompson5910 (jim_thompson5910):

so he'll have to eat at least 8 servings or more to make the avg 5 or more